If power of a bulb is 80 watt and wavelength of emitted radiation is 5000 Å, find number of photons emitted per second .

Grade:

1 Answers

Arun
5 years ago
Dear student
 
Energy of one photon = hc/lambda = 6.626* 10^-34 * 3 * 10^8 / 5000* 10^-10
 
= 3.976 * 10^-19 J
 
hence number of photons emitted per second = 80 / 3.976 *10^-19 = 20.121 *10^19